# Analytical Expression for Continuum-continuum Transition Amplitude of Hydrogen-like Atoms with Angular-momentum Dependence

### v1

@@ 2.1 Regular and irregular solutions to the Coulomb wave equation
where we used $h=\frac{t}{t+1}$. Compared to equation (49), we have $\alpha=a$, $\beta_{1}=\rho+1+n^{\prime}$, $\beta_{2}=-n^{\prime}$, $\gamma=\rho-b+a+2$, $z_{1}=1-\frac{\lambda}{Q}$, and $z_{2}=1-\frac{\lambda}{Q-\lambda^{\prime}}$. Therefore, we finally get
where in the last step we used the relation [52], equation (5.11.1):
With help of equations (41, 43, 44), we can calculate the multipole transition amplitudes between states expressed by the confluent hypergeometric functions of the first and the second kind. Therefore, equation (37) can be expressed as:  [^1] Here we make a special note that although $Q_{0}$ in equation (53) can be exactly zero, where the last two arguments in equation (51) are real and can be evaluated by analytic continuation when one of them has modulus greater than 1. However, due to the ambiguous phase of $(-1)$ that occurs upon analytic continuation, one of the branches of the parameters may become incorrect, which means that taking the complex conjugates of all the first four arguments in $F_{1}$ does not return the complex conjugate of the result, although from its definition one would expect so. Therefore, it is recommended to numerically confirm that the case of $Q_{0}=0$ is indeed the limit of $Q_{0}=0^{+}$.
One can find the similarity between our equations (41, 43, 44, 51) and the equation (3.1) in reference [50]. In fact, we can write
$=\int_{0}^{\infty}f_{l^{\prime}}(k^{\prime},r)~{}r\mathrm{e}^{-Q_{0}r}~{}u_{l}(k,r)~{}r^{2}\mathrm{d}r$
$=\int_{0}^{\infty}f_{l^{\prime}}(k^{\prime},r)~{}r\mathrm{e}^{-Q_{0}r}~{}g_{l}(k,r)~{}r^{2}\mathrm{d}r$
$+\mathrm{i}\int_{0}^{\infty}f_{l^{\prime}}(k^{\prime},r)~{}r\mathrm{e}^{-Q_{0}r}~{}f_{l}(k,r)~{}r^{2}\mathrm{d}r$ (54)
and one immediately sees that for the purpose of calculating the transition amplitude between two regular solutions to the Coulomb wave equation, also known as the Gordon’s integral, the result is given by the imaginary part of equation (54), which is also given in [48, 49, 50], while its real part corresponds to the transition amplitude between the regular and irregular solutions, where the regular and irregular solutions are defined to be real functions. We note that our formula has similarity to the general expression for the double- or multi-photon ionisation amplitude from bound states of hydrogen atoms given in references [53, 54], which use the Sturmaian expansion with modifications in the continuum.
2.3 Asymptotic formula for the CC transition amplitude with the centrifugal potential included
The purpose of this section is to demonstrate the mechanism with simpler mathematical expressions. Such approaches are usually available for describing the asymptotic behaviour at high kinetic energies, while we try to make the compromise that the formula applied at low kinetic energies should at least give a qualitative description, such that the main features are still captured. Let us consider the phase contribution of the centrifugal potential under the WKB approximation. The phase factor of the outgoing wave can be written as:
$S_{l}(k,r)=\int_{0}^{r}\sqrt{\frac{2}{\mathfrak{r}}-\frac{2b}{\mathfrak{r}^{2}}+k^{2}}~{}\mathrm{d}\mathfrak{r}$
$=\int_{0}^{r}k+\frac{\mathfrak{r}^{-1}}{k}-\frac{\mathfrak{r}^{-2}(2k^{2}b+1)}{2k^{3}}+\mathcal{O}(\mathfrak{r}^{-3})~{}\mathrm{d}\mathfrak{r}$
$=kr+\ln(2kr)/k+r^{-1}\frac{2k^{2}b+1}{2k^{3}}+\mathcal{O}(r^{-2})$ (55)
$rR_{k,l}(r)\approx\exp(\mathrm{i}S_{l}(k,r))$
$=\mathrm{e}^{\mathrm{i}kr}{(2kr)}^{\mathrm{i}/k}\exp(\mathrm{i}qr^{-1}+\mathcal{O}(r^{-2}))$
$=\mathrm{e}^{\mathrm{i}kr}{(2kr)}^{\mathrm{i}/k}\left(1+\mathrm{i}qr^{-1}+\mathcal{O}(r^{-2})\right)$ (56)
where $q=(2k^{2}b+1)/(2k^{3})$ scales with $k^{-3}$ for the $s$ ($l=0$) wave, but with $k^{-1}$ for waves with higher angular momenta. This explains why the phase contribution of the centrifugal potential should not be neglected, as $k^{-1}$ is a rather slow convergence under the experimental context (typically $E={10}^{1}\sim{10}^{2}~{}\mathrm{eV}$, $k=1\sim 3~{}\mathrm{a.u.}$). The dipole transition between two outgoing waves can thus be approximated as:
@@ 2.3 Asymptotic formula for the CC transition amplitude with the centrifugal potential included
$\mathcal{T}^{(\mathrm{P})}_{l\rightarrow l^{\prime}}(k,k^{\prime};r_{0})\approx-\frac{\pi}{2}N_{k}N_{k^{\prime}}\frac{{(2k)}^{\mathrm{i}/k}}{{(2k^{\prime})}^{\mathrm{i}/k^{\prime}}}\int_{r_{0}}^{\infty}\mathrm{e}^{-\mathrm{i}(k^{\prime}-k)r}~{}r^{\mathrm{i}/k-\mathrm{i}/k^{\prime}}\left[r+\mathrm{i}(q-q^{\prime})+\mathcal{O}(r^{-1})\right]~{}\mathrm{d}r$
$\approx-\frac{\pi}{2}N_{k}N_{k^{\prime}}\frac{{(2k)}^{\mathrm{i}/k}}{{(2k^{\prime})}^{\mathrm{i}/k^{\prime}}}\left[\frac{\Gamma(s,\Lambda r_{0})}{\Lambda^{s}}+\mathrm{i}(q-q^{\prime})\frac{\Gamma(s-1,\Lambda r_{0})}{\Lambda^{s-1}}\right]$ (57)
where $s=2+\mathrm{i}/k-\mathrm{i}/k^{\prime}$ and $\Lambda=\mathrm{i}(k^{\prime}-k)$. The factor of $1/2$ comes from the fact that the final state represented by the Kummer’s function contains the $k^{\prime}$- and $(-k^{\prime})$-parts, while the $(-k^{\prime})$-part practically vanishes due to the large imaginary part in the gamma function [55]. Here we have used the incomplete gamma function
$\Gamma(s,x)=\int_{x}^{\infty}t^{s-1}\mathrm{e}^{-t}~{}\mathrm{d}t$ (58)
We note that Dahlström et al. also reported another formula (equation (15)) that has the correction for the amplitude factor according to the WKB approximation:
$rR_{k,l}(r)\approx\exp(\mathrm{i}S_{l}(k,r)){k}^{1/2}{\left(2/r+k^{2}\right)}^{-1/4}$ (59)
where the amplitude factor is
${\left(\frac{2}{k^{2}r}+1\right)}^{-1/4}=1-\frac{1}{2k^{2}r}+\mathcal{O}(r^{-2}).$ (60)
Following the similar treatment, equation (57) is modified as
@@
$\mathcal{T}^{(\mathrm{P+A})}_{l\rightarrow l^{\prime}}(k,k^{\prime};r_{0})\approx-\frac{\pi}{2}N_{k}N_{k^{\prime}}\frac{{(2k)}^{\mathrm{i}/k}}{{(2k^{\prime})}^{\mathrm{i}/k^{\prime}}}\left[\frac{\Gamma(s,\Lambda r_{0})}{\Lambda^{s}}+(\mathrm{i}q-\mathrm{i}q^{\prime}-Q-Q^{\prime})\frac{\Gamma(s-1,\Lambda r_{0})}{\Lambda^{s-1}}\right]$ (61)
where $Q=1/(2k^{2})$. The superscripts $\mathrm{(P)}$ and $\mathrm{(P+A)}$ in equations (57, 61) refer to including the phase variation and the phase and amplitude variations, respectively. For the hydrogen atom at low kinetic energies, the phase computed by TDSE typically lies between equation (57) and equation (61) with $r_{0}=0$. Although the agreement may improve by choosing a suitable $r_{0}$ in a similar fashion as in the (P+A) model of reference [23], it is better to evaluate according to equation (53).
2.4 Numerical evaluation
3 Results and discussion
3.1 Continuum-continuum transition phase for the hydrogen atom
In order to test the accuracy of the analytical formula and its asymptotic approximations, we compare to the time-dependent Schrödinger equation (TDSE) results reported in [23, 36], which are currently considered to be the benchmark for the CC transition amplitudes. As shown in figure 3, our equation (53) perfectly agrees with the TDSE values, which manifests the validity of our approach. Regarding the asymptotic approximations, the (P) and (P+A) formulae given by Dahlström et al. are independent of the angular momentum $l^{\prime}$ (isotropic), and the CC phases of $p\rightarrow s$ and $p\rightarrow d$ lie between the two approximations. The modified asymptotic approximation (P) given by equation (57) reproduces the phase difference between the $p\rightarrow s$ and $p\rightarrow d$ channels, indicating that the effect of the centrifugal potential is indeed captured by the expansion, although they both overestimate the absolute value of the phase at low kinetic energies, as in the case of the isotropic formula. On the other hand, the modified asymptotic approximation (P+A) given by equation (61) show very little difference between $p\rightarrow s$ and $p\rightarrow d$, and both of them are well aligned with the corresponding isotropic formula. This is because at low kinetic energies the amplitude-variation effect overrides the phase-variation effect, while the former is independent of $l^{\prime}$ under the far-field expansion up to $\mathcal{O}(r^{-1})$.
Figure 3: The phase of the CC transition (with dressing field wavelength of 800 nm) for the hydrogen atoms from the intermediate state of $l=1$ to the final state of $l^{\prime}=1\mp 1$. The electron kinetic energy refers to the final state ($E^{\prime}$). The analytical formula (ana.) corresponds to (53); the $l$-dependent asymptotic approximation with phase correction (asym. (P), $r_{0}=0$) corresponds to (57); the $l$-dependent asymptotic approximation with phase and amplitude correction (asym. (P+A), $r_{0}=0$) corresponds to (61); the isotropic asymptotic approximation with phase correction (iso. asym. (P)) and with phase and amplitude correction (iso. asym. (P+A)) are taken from [23]; the TDSE results are taken from [36], which practically overlap with the SOPT calculation.
3.2 Fano’s propensity rule in continuum-continuum transition
Fano’s propensity rule in CC transition reported in references [34, 36] is tested using our equation (53) and its asymptotic approximations, as shown in figure 4. The analytical formula again shows excellent agreement with TDSE [23, 36], SOPT [60], and RPAE [34] for intermediate states with various angular momenta. We note that the modified asymptotic approximation shows very good agreement from about 15 eV, where the absorption pathway converges seemingly faster than the emission pathway. The (P) and (P+A) approximations ($r_{0}=0$) give very similar moduli ratios, as the amplitude variation effect is mostly cancelled in the ratios. The Fano’s propensity rule is entirely absent from the isotropic asymptotic formulae, since they do not have $l$-dependence. This indicates that the main effect from the centrifugal potential is well described by the expansion of the phase to the next order, which is a good approximation for $k\gtrsim 1~{}\mathrm{a.u.}$.
Figure 4: Moduli ratios for the absorption and emission pathways (with dressing field wavelength of 800 nm). The analytical formula (ana.) corresponds to (53) and the results are shown by the thick solid curves, while the results from the $l$-dependent asymptotic approximation (57,61) are shown by the thin dashed curves with the corresponding colours, where the results from the (P) and (P+A) formulae almost overlap, so only the ones from (P) are shown. The computational results (RPAE, SOPT, TDSE) are taken from [36], where the RPAE results originate from [34]. The angular momentum $l$ corresponds to the intermediate state, while the electron kinetic energy corresponds to the final state.
3.3 Continuum-continuum transition time delay
The continuum-continuum transition time delay defined by the finite difference is
where $E_{\pm}=E\pm E_{\mathrm{CC}}$ is the electron kinetic energy of the intermediate state, and $\phi$ is the phase of the transition amplitude. The results given by the analytical formula (53) are plotted in figure 5. The values for the $Ep\rightarrow E^{\prime}s$ and $Ep\rightarrow E^{\prime}d$ channels agree perfectly with the TDSE for hydrogen atoms. The (P) and (P+A) formulae reported by Dahlström et al. [23] and with the modification introduced in this work lie below and above the curves, respectively. In particular, the (P+A) formula bends to the opposite direction at lower kinetic energies, although it matches the TDSE better at higher kinetic energies. This again manifests the validity of our analytical formula. Interestingly, although the phases of different $l$-channels vary, as shown in figure 3, the phase differences between the emission and absorption pathways ($l\leq 3$) are very close, which has recently been reported by Busto et al. [21]. Hence, the key problem is to correctly determine the relative phase offset between channels and their moduli ratios, which can be accurately achieved by the analytical formula, as demonstrated in the previous sections.
Figure 5: The CC time delays obtained from the phases shown in figure 3. The TDSE results from hydrogen atoms are taken from [36]; the isotropic asymptotic approximation are taken from [23]; the other markers and curves are given by the analytical formula (53) labelled as “(intermediate angular momentum) $\rightarrow$ (final angular momentum)”. The electron kinetic energy corresponds to the final state.
3.4 Angle-dependent time delay for helium with circular-XUV-circular-IR RABBIT
Figure 6: Angle-dependent RABBIT phases and corresponding relative time delays of helium for co-rotating (co) and counter-rotating (ct) XUV and dressing IR (800 nm) with final-state kinetic energy of 3.3 eV (SB18). The phases and relative time delays are referenced to the co-rotating case with electron emitted at $90^{\circ}$. The experimental (exp.) and TDSE values are reported in [35]. The analytical (ana.) results are obtained from (53), including the calculation using unshifted and shifted center of expansion by 1 Å, shown as the solid and dashed lines, respectively. The dotted lines correspond to the values calculated by the formula proposed by Boll et al. [36]. The one-photon ionisation amplitude is calculated using ePolyScat [19, 20].
An experimental test for the angular-momentum-dependent continuum-continuum transition amplitude are angle-resolved time-delay measurements on atoms [61, 33, 62]. For example, using the (isotropic) asymptotic wavefunction for the intermediate state and the Kummer’s function for the final state [36], Boll et al. showed that this analytical formula with partial approximation matches well with the linear-XUV-linear-IR RABBIT experiments regarding the angle-dependent phase [37]. We note that due to the relatively large experimental uncertainties at emission angles close to $90^{\circ}$, where different theoretical models differ most, it is not obvious to decide which model is the most accurate, especially at lower electron kinetic energies. Here we compare the theories with a circular-XUV-circular-IR RABBIT experiment reported recently by the present authors [35]. The scheme for the generation and characterisation of the circular XUV attosecond pulse trains and the experimental setup for angle-resolved RABBIT have been described in our recent publications [35, 63, 64]. The benefit of the circular-XUV-circular-IR RABBIT scheme for helium is that the difference in the co- and counter-rotating XUV and IR provides additional information of the partial waves. For example, if the circular XUV excites one of the helium electrons from $1s$ to the $Ep_{1}$ intermediate state (the $z$-axis is defined as the light-propagation direction), in the continuum-continuum transition step, if the dressing field is co-rotating with the XUV, then the absorption pathway leads only to the $E^{\prime}d_{2}$ final state due to the restriction of the magnetic quantum number, while the emission pathway leads to the superposition of the $E^{\prime}s$ and $E^{\prime}d_{0}$ states; on the contrary, if the dressing field is counter-rotating, then the emission pathway yields the $E^{\prime}d_{2}$ final state, and the absorption pathway gives the superposition of the $E^{\prime}s$ and $E^{\prime}d_{0}$ states. Since there are only three partial waves contributing to a given sideband, their individual moduli and phases can be extracted by the global fitting of the two-dimensional interference pattern [65] with only six parameters. By comparing the co-rotating and counter-rotating cases, one can separate the Wigner part $1s\rightarrow Ep$ and the continuum-continuum part $Ep\rightarrow E^{\prime}s/d$ of the photoionsation time delays for each final state [35]. Theoretically, the angular dependence of the RABBIT phase is entirely absent from the isotropic asymptotic approximations, and the relative phase and modulus ratio between $Ep\rightarrow E^{\prime}s$ and $Ep\rightarrow E^{\prime}d$ play the key role, as explained in [34]. The theoretical and experimental results are compared in figure 6. We choose SB18 with relatively low electron kinetic energy (3.3 eV), where the asymptotic formulae have poor accuracy and are not shown in the figure. Our proposed analytical formula (53) shows remarkably good agreement with the TDSE and experimental values for both the co- and counter-rotating cases, which is expected from the level of accuracy that it has shown in describing the moduli and phases of the transition amplitudes between different $l^{\prime}$-states. The formula proposed by Boll et al. is known to have less accuracy at lower electron kinetic energies and agrees with the TDSE and experimental values semi-quantitatively. Considering the precision of the experimental values, we conclude that our analytical formula quantitatively reproduces the experimental results and has comparable accuracy with the TDSE computation, which outperforms the formula proposed by Boll et al. at low electron kinetic energies.
$E_{\rm k}$ (eV) $l=0$ $l=1$ $l=2$ $l=3$ $l>3$
3.3 0.0808 0.7484 0.1617 0.0088 0.0002
18.8 0.1891 0.2266 0.3970 0.1599 0.0274
In order to further check the sensitivity of our method regarding the shape of the potential, we performed calculations with helium one-photon photoionisation amplitudes with an expansion center displaced by 1 Å using ePolyScat [19, 20]. Since ePolyScat uses single-center expansions, when the atomic orbitals are displaced from the center of expansion, a series of (in principle infinite number of) partial waves are involved, in contrast to the centered helium atom, where the one-photon ionisation solely leads to the $p$-wave. The fractions (defined as the modulus square ratio) of partial waves of the shifted helium atom with electron kinetic energies of 3.3 eV and 18.8 eV are summarised in table 1. At 3.3 eV, $l=1$ ($p$-wave) is still dominating, with some contributions from $l=0$ and $l=2$ partial waves. As the electron kinetic energy increases to 18.8 eV, the distribution becomes more diffuse, with $l=2$ as the leading partial wave and comparable contributions from $l=0,1,3$. Although all the of observables at asymptotic distance from the origin should be insensitive to the shift of the atom, depending on the model, this may introduce artefacts. When the atom is displaced from the origin in the spherical coordinate, the potential is no longer central, and the unshifted Coulomb waves are no longer its eigenfunctions. On the other hand, one may argue that since the continuum-continuum transition involves the outgoing electron wavepackets that escape from the atom, the transition amplitude should be less sensitive to the short-range potential, and the asymptotic behaviour is the same. The results using the analytical formula for the lower kinetic energy (3.3 eV) are shown as the dashed lines in figure 6, where the phase difference compared to the unshifted case is only marginal. For the higher kinetic energy (18.8 eV), we compare the results using the analytical formula (53), the formula proposed by Boll et al. [36], and the modified asymptotic formula ((P), equation (57), $r_{0}=0$) in figure 7. Since the electron kinetic energy is relatively high, all three models show reasonable agreement with the experimental values. For the shifted helium, the time delay from the modified asymptotic formula is practically unchanged, since it is by its nature only sensitive to the asymptotic behaviour, whereas the formula proposed by Boll et al. shows a non-negligible phase difference, indicating that the time delay may be less accurate when the potential deviates from the perfect central potential. The analytical formula shows much smaller difference, which implies that although taking the confluent hypergeometric function of the second kind may deviate from the reality in a non-central potential field, e.g. the potential of a molecular cation, the calculated time delay is largely converged.
4 Conclusion
In conclusion, we have proposed an analytical formula for calculating the continuum-continuum transition amplitudes for the hydrogen-like atoms which uses the confluent hypergeometric functions of the first and second kinds as the final and intermediate states, respectively. We first derived the integral between the confluent hypergeometric functions of the first and second kinds, which can be expressed using the Appell’s $F_{1}$ function. We note that this expression is an extension to the Gordon’s integral, as the regular solution of the Coulomb wave equation can be expressed by the imaginary part of the irregular solution, or the linear combination of the two independent irregular solutions, and indeed one can find similarities between this expression and the reported expressions for the Gordon’s integral. Then we showed that this formula excellently reproduces the phase and the Fano’s propensity rule in the continuum-continuum transition, compared to the TDSE and other theoretical approaches. In addition, we modified the WKB asymptotic approximation by expanding the terms to $r^{-1}$, to which the centrifugal potential contributes. The Fano’s propensity rule and the main feature of the phases in continuum-continuum transitions can be nicely captured with this relatively simple expression that involves only the gamma function. Finally, we used our method to calculate the relative RABBIT phases between the co- and counter-rotating XUV and IR at different electron emission angles. Even at kinetic energy as low as 3.3 eV, our proposed analytical formula quantitatively agrees with both the TDSE calculation and the experimental values. We also checked the case with the shifted expansion center for calculating the spherical harmonics, and the results indicates that our method is not very sensitive to the shape of the short-range potential field, which opens the possibility of adapting our method to molecules.
Acknowledgements
We thank David Busto and Mathieu Gisselbrecht for fruitful discussions. This work was funded by ETH Zürich. J.-B. Ji acknowledges the funding from the ETH grant 41-20-2. M. H.’s work was funded by the European Union’s Horizon 2020 research and innovation program under Marie Skłodowska-Curie agreement grant No. 801459, FP-RESOMUS.

### v2

@@ Appendix A Derivation for the radial integral
where we used $h=\frac{t}{t+1}$. Compared to equation (65), we have $\alpha=a$, $\beta_{1}=\rho+1+n^{\prime}$, $\beta_{2}=-n^{\prime}$, $\gamma=\rho-b+a+2$, $z_{1}=1-\frac{\lambda}{Q}$, and $z_{2}=1-\frac{\lambda}{Q-\lambda^{\prime}}$. Therefore, we finally get equation (44) in the main text.
Appendix B Formula for arbitrary $Z>0$
For the general case of $Z>0$ instead of $Z=1$, equation (20) becomes
$f_{l}(k,r;Z)=|C_{l}(k;Z)|~{}r^{l}~{}\exp\left({\mathrm{i}kr}\right)\Phi\left(l+1-\mathrm{i}Z/k,2l+2,-2\mathrm{i}kr\right)$ (67)
where
$C_{l}(k;Z)={k}^{l+1}~{}{2}^{l}~{}\mathrm{e}^{{\pi Z}/{(2k)}}~{}\frac{\Gamma\left(l+1+\mathrm{i}Z/k\right)}{(2l+1)!}~{}.$ (68)
Similarly, equation (22) becomes
$u_{l}(k,r;Z)=B_{l}(k;Z)~{}|C_{l}(k;Z)|~{}r^{l}~{}\exp\left({\mathrm{i}kr}\right)U\left(l+1-\mathrm{i}Z/k,2l+2,-2\mathrm{i}kr\right)$ (69)
with
$B_{l}(k;Z)=-2\mathrm{i}\mathrm{e}^{-{\pi Z}/{k}}{(-1)}^{l}\frac{(2l+1)!}{\Gamma\left(l+1+\mathrm{i}Z/k\right)}~{}.$ (70)
Through the same derivation, the radial integral (46) can be generalised into:
$\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime};Z)=-\pi N_{k}N_{k^{\prime}}\mathrm{i}^{l-l^{\prime}-1}B_{l}(k;Z)C_{l}(k;Z)C_{l^{\prime}}^{*}(k^{\prime};Z)$ (71)
@@ Appendix B Formula for arbitrary $Z>0$
×$\times\lim_{Q_{0}\rightarrow 0^{+}}\mathcal{J}^{l-l^{\prime}+2}_{2l^{\prime}+1}\left(l+1-\frac{\mathrm{i}Z}{k},2l+2,-2\mathrm{i}k;-\left(l^{\prime}+1-\frac{\mathrm{i}Z}{k^{\prime}}\right),-2\mathrm{i}k^{\prime};Q_{0}-\mathrm{i}(k+k^{\prime})\right)$
and its value can be obtained using equations (40) and (44).
Appendix C Formula for multipole transitions
If the transition term is $r^{\mathcal{N}}$ instead of $r$ in equation (36), where $\mathcal{N}$ is a positive integer, then equation (71) can be further generalised as:
$\mathcal{T}^{\mathcal{N}}_{l\rightarrow l^{\prime}}(k,k^{\prime};Z)=-\pi N_{k}N_{k^{\prime}}\mathrm{i}^{l-l^{\prime}-1}B_{l}(k;Z)C_{l}(k;Z)C_{l^{\prime}}^{*}(k^{\prime};Z)$ (72)
@@ Appendix C Formula for multipole transitions
×$\times\lim_{Q_{0}\rightarrow 0^{+}}\mathcal{J}^{l-l^{\prime}+1+\mathcal{N}}_{2l^{\prime}+1}\left(l+1-\frac{\mathrm{i}Z}{k},2l+2,-2\mathrm{i}k;-\left(l^{\prime}+1-\frac{\mathrm{i}Z}{k^{\prime}}\right),-2\mathrm{i}k^{\prime};Q_{0}-\mathrm{i}(k+k^{\prime})\right)$
and its value can also be obtained using equations (40) and (44). The requirement of $l-l^{\prime}+1+\mathcal{N}\geq 0$ is guaranteed by the selection rule that $|l-l^{\prime}|\leq\mathcal{N}$.
Appendix D Explanation of Cooper-like minima
The Cooper-like minimum in the absorption pathway of the $l\rightarrow l-1$ channel of CC transition originates from the transition amplitude between the irregular part of the intermediate-state wavefunction and the final-state wavefunction. Let us consider the asymptotic behaviour of the function $u_{l}(k,r)$ in expression (26). The irregular ($g_{l}(k,r)$) and the regular ($f_{l}(k,r)$) parts behave like the cosine and sine functions in the far field, respectively. Due to the factor ${(-\mathrm{i})}^{l}$, the wave of two neighboring $l$s will have phase shift of $\pi/2$; i.e. the sine and cosine are swapped. Following that the intermediate state has both the regular and irregular parts while the final state has only the regular part, the regular-regular transition can be roughly expressed as
$\int_{0}^{\infty}\sin(kr)\cos(k^{\prime}r)r\mathrm{d}r=\frac{1}{2}\int_{0}^{\infty}\sin((k+k^{\prime})r)+\sin((k-k^{\prime})r)\mathrm{d}r$ (73)
while the irregular-regular transition reads:
$\int_{0}^{\infty}\cos(kr)\cos(k^{\prime}r)r\mathrm{d}r=\frac{1}{2}\int_{0}^{\infty}\cos((k+k^{\prime})r)+\cos((k-k^{\prime})r)\mathrm{d}r~{}.$ (74)
These integrals can be evaluated using the following equations:
$\int_{0}^{\infty}\sin(\mathfrak{K}r)\mathrm{e}^{-\lambda r}~{}r~{}\mathrm{d}r=\frac{1}{\mathfrak{K}^{2}}\frac{2\lambda/\mathfrak{K}}{{\left({(\lambda/\mathfrak{K})}^{2}+1\right)}^{2}}$ (75)
$\int_{0}^{\infty}\cos(\mathfrak{K}r)\mathrm{e}^{-\lambda r}~{}r~{}\mathrm{d}r=\frac{1}{\mathfrak{K}^{2}}\frac{{(\lambda/\mathfrak{K})}^{2}-1}{{\left({(\lambda/\mathfrak{K})}^{2}+1\right)}^{2}}$ (76)
where $\mathfrak{K}=k\pm k^{\prime}$. By taking the limit $\lambda\rightarrow 0^{+}$, (75) and (76) approach 0 and $-1/\mathfrak{K}^{2}$, respectively, and in the latter integral $\mathfrak{K}=k-k^{\prime}$ contributes significantly more than $\mathfrak{K}=k+k^{\prime}$, provided that the dressing-field photon energy is smaller than the final-state electron kinetic energy. Therefore, we have the integrals
$\mathcal{I}^{\mathrm{reg\rightarrow reg}}_{l\rightarrow l\pm 1}\sim 0,$ (77)
and
$\mathcal{I}^{\mathrm{reg\rightarrow irr}}_{l\rightarrow l\pm 1}\propto-\frac{1}{2{\Delta k}^{2}},$ (78)
where $\Delta k\equiv k-k^{\prime}=(k^{2}-{k^{\prime}}^{2})/(k+k^{\prime})=\Delta E/\bar{k}\approx\Delta E/{(2\bar{E})}^{1/2}$, $\Delta E\equiv E-E^{\prime}=\mp E_{\mathrm{CC}}$ for the absorption ()(-)( - ) and emission (+)(+)( + ) pathways, $\bar{k}\equiv(k+k^{\prime})/2$ stands for the averaged momentum of the intermediate and final states, and $\bar{E}\equiv(E+E^{\prime})/2$ is the averaged electron kinetic energy. On the other hand, however, the assumption that the wavefunctions have the oscillating feature is only true, according to the WKB approximation, when the classical local kinetic energy is positive, which means
$k^{2}+\frac{2}{r}-\frac{l(l+1)}{r^{2}}=0$ (79)
according to equation (19) under $Z=1$. When $l$ is large and $r$ is relatively small, the classical turning point is $\sim\sqrt{l(l+1)}/k\sim l/k$, which leads to the plateau for the $l\rightarrow l-1$ transition shown in figure 2 (b) (red curve) between 0 and the classical turning point, i.e. inside the centrifugal barrier, where the integral in this particular part is
$\mathcal{I}^{\mathrm{barrier}}_{l\rightarrow l-1}\propto\int_{0}^{l/k}r\mathrm{d}r=\frac{l^{2}}{2k^{2}}.$ (80)
Note that the integral in (76) also has contribution inside the barrier, but there is a factor of $1/2$ so it contributes less than the exact formula, and overall the exact value is added by a positive component $\zeta^{2}l^{2}/(2k^{2})$, where $\zeta>0$. Using (78), the total integral crosses zero when $\zeta^{2}l^{2}/(2k^{2})=1/({2\Delta k}^{2})={\bar{k}}^{2}/(2{\Delta E}^{2})$, namely $k^{2}\approx\zeta lE_{\mathrm{CC}}$, which gives the observed linear scaling of the positions of the minima. On the contrary, as shown in figure 2 (c), for the $l\rightarrow l+1$ transition, the product of the irregular part of the intermediate state and the final state near the origin approaches zero, which leads to a negative term added to the overall integral and avoids the zero point.
This can also be semi-quantitatively derived from the modified asymptotic formula (50). Letting $Z=1$ and assuming $k$ is relatively large, we use the rough approximation of $\Gamma(2+\mathrm{i}/k-\mathrm{i}/k^{\prime})\approx\Gamma(2)=1$ and $\Gamma(1+\mathrm{i}/k-\mathrm{i}/k^{\prime})\approx\Gamma(1)=1$, and $|{[\mathrm{i}(k^{\prime}-k)]}^{2+\mathrm{i}/k-\mathrm{i}/k^{\prime}}|\approx{\Delta k}^{2}$, so the zero of $|\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})|$ is reached when $01+(k-k^{\prime})(q-q^{\prime})=$ 0 in formula (50). For large $l$, $q\approx b/k=l(l+1)/(2k)$ and $q^{\prime}\approx b^{\prime}/k^{\prime}=l(l-1)/(2k^{\prime})$, where we have used $l^{\prime}=l-1$. Therefore, $2kk^{\prime}(q-q^{\prime})=l^{2}(k^{\prime}-k)+l(k^{\prime}+k)$, and the condition for the zero becomes:
$0=1-\frac{E_{\mathrm{CC}}}{\bar{k}}\left(\frac{E_{\mathrm{CC}}}{\bar{k}}\frac{l^{2}}{2kk^{\prime}}+\frac{l\bar{k}}{kk^{\prime}}\right)=1-\frac{lE_{\mathrm{CC}}}{kk^{\prime}}-\frac{l^{2}{E_{\mathrm{CC}}}^{2}}{2{\bar{k}}^{2}kk^{\prime}}\approx 1-\frac{lE_{\mathrm{CC}}}{{\bar{k}}^{2}}-\frac{1}{2}\frac{l^{2}{E_{\mathrm{CC}}}^{2}}{{\bar{k}}^{4}}.$ (81)
Solving this equation yields
$\bar{E}={\bar{k}}^{2}/2=\left(\frac{1+\sqrt{3}}{4}\right)lE_{\mathrm{CC}}.$ (82)
Although this coefficient ($\sim 0.7$) is smaller than the observation ($\sim 1.0$), considering the level of approximation, this has reflected the main feature of the behaviour of the transition amplitude between the irregular part of the intermediate state with $l$ and the final state with $l-1$ by absorbing one photon from the dressing field.